Hello, Can someone please help me with this problem.
Given that P(A) = 0.5 and P(B) = 0.2, and P(A and B) = 0.14, determine P(B|A). (Points : 3)
0.10
0.28
0.70
0.20
Thank you, I been out of school for 15 years and I confused when it come to these type of problems I want to say it 10 but I not sure.

Answer by stanbon(75887)

Given that P(A) = 0.5 and P(B) = 0.2, and P(A and B) = 0.14, determine P(B|A). (Points : 3)
0.10
0.28
0.70
0.20
-----
P(B|A) = P(A+B)/P(A) = 0.14/0.50 = 0.28
=====================
Cheers,
Stan H.